OpenStudy (anonymous):

HELP?? If 5.65 grams of zinc metal react with 21.6 grams of silver nitrate, how many grams of silver metal can be formed and how many grams of the excess reactant will be left over when the reaction is complete? Show all of your work. unbalanced equation: Zn + AgNO3 Zn(NO3)2 + Ag

OpenStudy (chmvijay):

can you write the balanced equation and tell me the molecualr weight of AgNO3 and Zn(NO3)2 and atomic weight of Ag and Zn

OpenStudy (anonymous):

Zn + 2(AgNO3) = Zn(NO3)2 + 2Ag

OpenStudy (australopithecus):

can you balance the equation? Once you have balanced it, you need to convert grams to moles for both compounds (you know how to do that right?) Then figure out what the limiting reagent (the reactant that is used up first, and thus dictates when the reaction is complete). To do this you simply see how many moles of product both reactants will give, the reactant that gives the least amount (smallest amount of moles) is your limiting reagent. then simply use the moles of your limiting reagent to figure out how much product is created, and how much of the other reactant is used. Subtract the amount of the non limiting reactant used from total amount of non limiting reagent used to figure out how much excess reactant is left over

OpenStudy (australopithecus):

Any questions?

OpenStudy (anonymous):

Yes): I don't know how to do it at all

OpenStudy (chmvijay):

ok fine can you do this find out molecualr weight of AgNO3 and Zn(NO3)2 and atomic weight of Ag and Zn

OpenStudy (australopithecus):

\[Moles = \frac{Grams}{MolecularMass}\] Molecular Mass is a constant you get from your periodic table, grams are given in this question To find molecular mass of CO2 for example The molecular mass of Carbon = 12g/mol The molecular mass of Oxygen = 16g/mol 12g/mol + 16g/mol + 16g/mol = 44g/mol therefore, CO2 has a molecular mass of 44g/mol

OpenStudy (australopithecus):

can you find the moles of both reactants now?

OpenStudy (australopithecus):

If you dont follow I can give you another example

OpenStudy (australopithecus):

the unit for moles is mol btw

OpenStudy (anonymous):

I kind of understand your example. So i do that for AgNO3?

OpenStudy (australopithecus):

Well CO2 has One Carbon atom and Two Oxygen atoms so we add up all the individual atoms molecular weight to find the molecular weight of the compound so, The molecular mass of Carbon + The molecular Mass of Oxygen + the molecular Mass of oxygen = The molecular mass of CO2

OpenStudy (australopithecus):

yes you do that, then you use the formula for moles I provided (be sure to memorize it) to find the moles of each of the reactants

OpenStudy (australopithecus):

Another example, NO I have one nitrogen and one oxygen atom, molecular mass of nitrogen = 14g/mol molecular mass of oxygen = 16g/mol 14g/mol + 16g/mol = 30g/mol therefore the molecular mass of NO is 30g/mol if I had 30 grams of it and I wanted to convert to moles, \[\frac{30g}{30g/mol} = 1 mol\] therefore, I have 1 mol of NO for every 30g of NO

OpenStudy (anonymous):

Well i understand your examples i just don't know how to do that for my problem.

OpenStudy (australopithecus):

get a periodic table and look up the molecular mass of each atom and find the total molecular mass of each compound then use the ratio I provided to find the moles

OpenStudy (australopithecus):

you need to find moles of both of your reactants, just follow the steps I did and you will be closer to solving your problem.

OpenStudy (australopithecus):

I will explain how to solve your problem using another example

OpenStudy (australopithecus):

well how to solve it once you find the moles of both reactants

OpenStudy (australopithecus):

You can post your work I dont mind checking it

OpenStudy (anonymous):

So Zn= 65.39 ...

OpenStudy (australopithecus):

Correct, but always include units, g/mol you will lose marks and its very essential in industry

OpenStudy (australopithecus):

Now find the moles of Zn(NO3)2 you have

OpenStudy (anonymous):

This is hopeless. Im so confused. Im sorry in waisting your time):

OpenStudy (australopithecus):

If I had the reaction: 2NO + O2 → 2 NO2 I reacted 30g of NO and 32grams of O2 and I wanted to find the amount of each reactant left over and the amount of product formed I would first convert over to moles 30g/30g/mol = 1mol of NO 32g/32g/mol = 1 mol of O2 Then I use the ratios from the balanced equation for every 2 moles of NO I use I get 2 moles of NO2 so, \[\frac{1mol_{NO}}{2}*2 = 1mol_{NO_2}\] for every 1 mole of O2 I get 2 Moles of NO2 so, \[\frac{1mol_{O_2}}{1}*2 = 2mol_{NO_2}\] NO is the limiting reagent because it takes more moles of NO to make 1mol of NO2 than it does O2. So we know NO is going to be used up before O2 will be in this reaction, therefore we can use it to determine the total amount of product formed and the amount of O2 left over once the reaction is complete assuming all the limiting (NO) is consumed in the reaction. So we just apply the ratios again from our equation, we have 1mol of NO (the 30g) for every 2 moles of NO consumed we get 2 moles of NO from earlier we saw that we get 1 mole of product from 1 mole of NO, to find how much O2 was consumed in the reaction we simply back track from the product so, \[\frac{1mol_{NO_2}}{2}*1 = 0.5mol_{O_2}\] thus 0.5moles of O2 are consumed in the reaction, To find the amount of left over O2 we simply use the formula Left Over Reagent = Total Reagent utilized - Amount of Reagent Reacted so in our case, 0.5mol O2 Left Over = 1mol O2 - 0.5mol O2 reacted

OpenStudy (australopithecus):

You can do this stuff it isn't that hard, you just need practice :) If you have any questions I'm here to help

OpenStudy (australopithecus):

You aren't wasting my time

OpenStudy (anonymous):

Okay. sorry I'm replying a lot I'm just trying to process what your explaining to me

OpenStudy (australopithecus):

So you need to find the molecular mass of, AgNO3 It consists of 1 Ag atom 1 N atom 3 O atoms Find the molecular mass of Oxygen, Nitrogen and Silver then use the following formula 1*Molecular Mass Ag + 1*Molecular Mass N + 3*Molecular Mass O

OpenStudy (australopithecus):

Yeah its fine take your time, I think I was pretty detailed in my explanation so I dont really have anything more to say ha.

OpenStudy (anonymous):

107.86Ag + 14.006N + 15.99Ox3=47.97 what do i do now?

OpenStudy (australopithecus):

Check out my example, :) dont get confused about the example I did having 1 product and yours having 2

OpenStudy (chmvijay):

@Australopithecus go ahead very nice :)

OpenStudy (australopithecus):

you can pick either of your products to determine what the limiting reagent is

OpenStudy (australopithecus):

in your reaction

OpenStudy (australopithecus):

The molecular mass you got for AgNO3 is wrong btw but I'm sure you will see that right away

OpenStudy (anonymous):

Yeah I'm working it out on a scratch piece of paper now. I'll let you know if i run into any more trouble. Thank you!!

OpenStudy (australopithecus):

Ok :) No problem